OpenStudy (anonymous):

Given w=f(x,y) x=rcos(theta) y=rsin(theta) a) compute 2w/2x and 2w/2y b)Prove that...

OpenStudy (anonymous):

\[(\frac{ 2w }{ 2x })^2+(\frac{ 2w }{ 2y })^2=(\frac{ 2w }{ 2r })^2+(\frac{ 1 }{ r^2 })(\frac{ 2w }{ 2theta })^2\]

ganeshie8 (ganeshie8):

do u mean partials ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

you may use \(\large \color{Red}{\text{\partial}}\) for showing \(\large \partial \)

OpenStudy (anonymous):

I don't know how to start the problem though

ganeshie8 (ganeshie8):

start by finding each partial

ganeshie8 (ganeshie8):

\(\large \frac{\partial w}{\partial r} = ? \) \(\large \frac{\partial w}{\partial \theta} = ? \)

ganeshie8 (ganeshie8):

\(\large \frac{\partial w}{\partial r} = \frac{\partial w}{\partial x}\frac{\partial x}{\partial r} + \frac{\partial w}{\partial y}\frac{\partial y}{\partial r} \) \(\large \qquad= \frac{\partial w}{\partial x}\cos\theta + \frac{\partial w}{\partial y}\sin\theta \)

ganeshie8 (ganeshie8):

similarly find \(\large \frac{\partial w}{\partial \theta} \), and plug them in ur right hand side of the equation

ganeshie8 (ganeshie8):

and try to simplify it to get the left hand side

OpenStudy (anonymous):

Is it \[\frac{ 2w }{ 2 \theta }=\frac{ 2w }{ 2x }\frac{ 2x }{ 2 \theta }+\frac{ 2w }{ 2y }\frac{ 2y }{ 2 \theta }\]?

ganeshie8 (ganeshie8):

thats right, and since \(\large x = r \cos \theta\) and \(\large y = r\sin \theta\) just evaluate these partials and plug in above...

OpenStudy (anonymous):

Is it \[\frac{ 2w }{ 2 \theta }=\frac{ 2w }{ 2x }(-rsin \theta)+\frac{ 2w }{ 2y }(rcos \theta)\]?

ganeshie8 (ganeshie8):

looks perfect !

OpenStudy (anonymous):

So then do I solve and see if I get the equation from the one I first posted?

OpenStudy (anonymous):

To prove it?

ganeshie8 (ganeshie8):

just plug these on right hand side

OpenStudy (anonymous):

Could you set it up, I'm still kind of confused

ganeshie8 (ganeshie8):

Right hand side of the equation you wanted to prove : \(\large \left(\frac{\partial w}{\partial r} \right)^2 + \frac{1}{r^2}\left(\frac{\partial w}{\partial \theta} \right)^2 \)

OpenStudy (anonymous):

ganesh=wrong... don't listen hehe

OpenStudy (anonymous):

So plug the whole answers into 2w/2r and 2w/2theta to get the left hand side?

ganeshie8 (ganeshie8):

we have a tough guest here :/

ganeshie8 (ganeshie8):

yes !

OpenStudy (anonymous):

ok ill give it a try

ganeshie8 (ganeshie8):

right after plugging in, most terms cancel out and take u to left hand side should be easy...

ganeshie8 (ganeshie8):

good luck !

OpenStudy (anonymous):

One question, how do I find 2w/2x and 2w/2y themselves? Or would 2w/2r and 2w/2theta be good enough?

ganeshie8 (ganeshie8):

leave them as they are.. you dont have them on ur right hand side, eh ?

ganeshie8 (ganeshie8):

Right hand side : \(\large \left(\frac{\partial w}{\partial r} \right)^2 + \frac{1}{r^2}\left(\frac{\partial w}{\partial \theta} \right)^2 \) \(\large \left(\frac{\partial w}{\partial x}\cos\theta + \frac{\partial w}{\partial y}\sin\theta \right)^2 + \frac{1}{r^2}\left(\frac{\partial w}{\partial x}(-r\sin\theta) + \frac{\partial w}{\partial y}(r\cos\theta)\right)^2 \)

OpenStudy (anonymous):

That's what I got so far, I'm just squaring it right now, is it suppose to be long?

ganeshie8 (ganeshie8):

nope, next step is ur answer

ganeshie8 (ganeshie8):

Right hand side : \(\large \left(\frac{\partial w}{\partial r} \right)^2 + \frac{1}{r^2}\left(\frac{\partial w}{\partial \theta} \right)^2 \) \(\large \left(\frac{\partial w}{\partial x}\cos\theta + \frac{\partial w}{\partial y}\sin\theta \right)^2 + \frac{1}{r^2}\left(\frac{\partial w}{\partial x}(-r\sin\theta) + \frac{\partial w}{\partial y}(r\cos\theta)\right)^2 \) \(\large \left(\frac{\partial w}{\partial x}\cos\theta + \frac{\partial w}{\partial y}\sin\theta \right)^2 + \frac{1}{r^2}\times r^2\left(\frac{\partial w}{\partial x}(-\sin\theta) + \frac{\partial w}{\partial y}(\cos\theta)\right)^2 \)

ganeshie8 (ganeshie8):

^^r^2 cancels out

OpenStudy (anonymous):

where did you get the r^2 from?

OpenStudy (anonymous):

Oh wait nvm you factored the r^2 out

ganeshie8 (ganeshie8):

Right hand side : \(\large \left(\frac{\partial w}{\partial r} \right)^2 + \frac{1}{r^2}\left(\frac{\partial w}{\partial \theta} \right)^2 \) \(\large \left(\frac{\partial w}{\partial x}\cos\theta + \frac{\partial w}{\partial y}\sin\theta \right)^2 + \frac{1}{r^2}\left(\frac{\partial w}{\partial x}(-r\sin\theta) + \frac{\partial w}{\partial y}(r\cos\theta)\right)^2 \) \(\large \left(\frac{\partial w}{\partial x}\cos\theta + \frac{\partial w}{\partial y}\sin\theta \right)^2 + \frac{1}{r^2}\times r^2\left(\frac{\partial w}{\partial x}(-\sin\theta) + \frac{\partial w}{\partial y}(\cos\theta)\right)^2 \) \(\large \left(\frac{\partial w}{\partial x}\cos\theta + \frac{\partial w}{\partial y}\sin\theta \right)^2 +\left(\frac{\partial w}{\partial x}(-\sin\theta) + \frac{\partial w}{\partial y}(\cos\theta)\right)^2 \) \(\left(\frac{\partial w}{\partial x} \right)^2\cos^2\theta + \left(\frac{\partial w}{\partial y} \right)^2\sin^2\theta + \left(\frac{\partial w}{\partial x} \right)^2\sin^2\theta + \left(\frac{\partial w}{\partial y} \right)^2\cos^2\theta \)

ganeshie8 (ganeshie8):

2ab terms cancel away because of opposite signs.

ganeshie8 (ganeshie8):

you're left only with the square terms... see if you're okay so far

OpenStudy (anonymous):

Yep I think I got it right so far, do I factor out the sin and cos?

ganeshie8 (ganeshie8):

factor the partials instead so that we can use sin^2 + cos^2 = 1

ganeshie8 (ganeshie8):

group 1st and 3rd terms

OpenStudy (anonymous):

\[(\frac{ 2w }{ 2x })^2(\cos^2 \theta+\sin^2 \theta)+(\frac{ 2w }{ 2y })^2(\cos^2 \theta+\sin^2 \theta)\]Right?

ganeshie8 (ganeshie8):

yes !

OpenStudy (anonymous):

then replace cos^2+sin^2 to 1 to get the final answer. Right?

OpenStudy (anonymous):

Thanks!

ganeshie8 (ganeshie8):

u wlc :)